OpenStudy (anonymous):

Find the radius of a cylinder with a volume of 188.5 feet³ and height of 15 feet.

OpenStudy (anonymous):

the formula for the volume, is V = (pi)r^2(h) so if the volume is 188.5; and the height is 15; we enter that information in the formula 188.5 = (pi)r^2(15) we pass the 15 to the other side, since its multiplying it; we divide both sides by it 188.5/15 = (pi)r^2 15/15 12.5666 = (pi)r^2 (1) ------ times one, is the same thing 12.5666 = (pi)r^2 now we pass the (pi) = 3.14 or 22/7; in this case we can use 3.14; which it will end as a approximation 12.5666/ 3.14 = r^2 4.0021 = r^2 --- approximately 4 4 = r^2 to figure that part out we find the square root of each *sqrt(4 = r 2 = r r = 2 so the radius will equal to 2